FastBits Electronic Company Sdn. Bhd. is evaluating new precision inspection devices to help verify package quality. The manager has obtained the following bids from four companies. All devices have a life of five years and a minimum attractive rate of return of 5% The alternatives are mutually exclusive Description Company A Company B Company C Company D Initial Cost (RM) 360000 116000 440000 200000 Annual Costs (RM) 900 Net Cash Flows (RM)90000 32480 110000 46200 IRR Determine the annual benefits of the devices from all four companies. Company A Company E CompanyC Company D Device from which company has the highest annual benefit? FastBits should reject the bid from which company based on the given individual IRR? Using incremental internal rate of return analysis, from which company, if any, should the manager purchase the new precision inspection device? Use trial and error method with 5% and 11% interest rates. Understood? (Y/N) Step 1- Eliminate Company Step 2 Rank Company from no 1-2-3 Step 4 - Incremental IRR first comparison Step 5- Remove Company from selection Repeat Step 4 Incremental IRR 2nd comparison Step 5 Choose Company Demonstrate that the same company selection would be made with proper application of the Present Worth (PW) method PW Company A PW Company B PW Company C PW Company D Thus, choose Company 12000 23000 9000 12.4% 7.9% 5% Format 50800 Format: 42820 Format: 485000 Format: 58500 Format: A Format: A Format: A Format: A Format xx-x Format: 7.3 Format: A Format: 5.8 Format A Format:96828 Format 67899 Format: 84976 Format : 72 Format: A
Answer #1

Annual Benefits

Company A : Net Cash Flows for 5 years is 90000 so annually it comes to 90000/5=18000

Company B : Net Cash Flows for 5 years is 32480 so annually it comes to 32480/5= 6496

Company C : Net Cash Flows for 5 years is 110000 so annually it comes to 110000/5=22000

Company D: Net Cash Flows for 5 years is 46200 so annually it comes to 46200/5=9240

Company C is providing Device with highest annual benefit of 22000

FastBits should reject the Bid of Company D as it offers lowest IRR

Step 1: Eliminate Company : Company D

Step 2: Rank Company 1-2-3

Rank 1: Company B; Rank 2 : Company A; Rank 3: Company C

Step 4: Incremental IRR Company B> Company A> Company C